Status Report

NASA Space Station On-Orbit Status 2 Jan 2004

By SpaceRef Editor
January 2, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 2 Jan 2004

All ISS systems continue to function nominally, except as noted previously or below.

Continuing the current round of monthly Russian Segment (RS) air ventilation systems preventive maintenance, FE Alexander Kaleri worked in the FGB for a cleanup of the protective mesh screens of the TsV ventilation fans (last time done: 12/12).

CDR/SO Michael Foale reviewed today’s planned CBOSS-FDI (Cellular Biotechnology Support Systems-Fluid Dynamics Investigation) payload procedure, then completed the particle injection and pumping on TCM (tissue culture module) A4 with a 3mL syringe.

Foale conducted another round of temperature readings in the Lab module, using the “Pilobolus” scopemeter with its temperature probe attachment for several data acquisitions during the day.

After Mike’s “ghosting” of PCS HDD #6095 (portable computer system/hard-disk drive) on 12/18/03, Kaleri today “cloned” two more HDDs (#6029, #6063) with the current 12A “ghost” image load, to add two functioning PCS to the working laptops (now five).  [Two HDDs remain as spares (#6058, #6054).  They can be prepared in multiple sessions, with each HDD cloning requiring one hour execute time.]

The FE performed the regular once-a-week maintenance reboot on all operational PCS laptops.  He also rebooted the OCA comm router laptop (every two weeks).

After the reconfiguration of EXPRESS Rack #4 (ER4) on 12/31, Mike Foale today performed POSSUM photo documentation on the rack.  [POSSUM is a payload photo activity that obtains electronic still shots of any subrack and locker payload that is moved or reconfigured.]

Mike then worked on ER1, performing a manual reboot of its RIC (rack interface controller) laptop.

Foale also completed today’s scheduled checkout and cleaning of the filters of the SAMS (space acceleration measurement system) hardware in both RTS (remote triaxial sensor) drawers and the ICU (interim control unit) drawer, using the vacuum cleaner.  [For the cleaning, the ground deactivated the ICU and both RTSs (D1, D2).  Rack switches and drawer power & fan switches were left on.]

In further preparation for the upcoming runs of the CGBA (Commercial Generic Bioprocessing Apparatus) experiment, which uses the reconfigured ER4, the CDR conducted an on-board training (OBT) review of the planned payload operations.

Mike also transferred data files from the physical exercise equipment (TVIS and RED) to the medical equipment computer (MEC).  Subsequently, he completed the periodic transfer of accumulated data files from the CMS HRM (crew medical systems/heart rate monitor) to the MEC, then erased them on the HRM.

Sasha performed the daily routine maintenance of the SOZh life support system (which includes ASU toilet facilities) and prepared the daily IMS (inventory management system) delta file for automatic export/import to update the database.  Mike conducted the regular routine status checkup of running ISS-8 payloads.

The CDR used the CDMK (carbon dioxide monitoring kit) to take the periodic reading of the cabin air’s current CO2 partial pressure, for calldown to the ground (along with the battery status) for use in trending analyses.  The CDMK was then restowed at the SM Central Post.

At 9:02am EST, Mike Foale set up and conducted a 10-min. ham radio session with students at Gilmour Academy in Gates Mill, OH.  [Gilmour Academy, located in Gates Mills near Cleveland, OH, has an amateur radio station, and today’s participating students, whose questions were uplinked beforehand, came from grades 4-12.]

Later, at 10:30am, Mike downlinked a non-interactive PAO video program with a view of the cabin interior and his thoughts on life and work on the station as well as a look at coming events.

Both crewmembers worked out with their daily 2.5-h program of physical exercise, on TVIS treadmill, RED expander, and VELO cycle with load trainer.  Mike used today’s TVIS session also to repair his TVIS harness, on which the front part of the shoulder strap had come off of the waist harness yesterday.

After yesterday’s successful attitude maneuver, ISS is now in earth-“fixed” LVLH (local vertical/local horizontal), to remain there until 1/8.

For today’s CEO (crew earth observations) sessions, the ground uplinked a list of target cities as potential photographic opportunities at the crew’s discretion.

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 8:00am EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On.  Vozdukh CO2 scrubber is On (in Manual Mode 3).  U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode).  TCCS (trace contaminant control subsystem) is operating.  MCA (major constituents analyzer) is off (in Life Extending Mode).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On; SKV-2 is Off (repair incomplete).

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Autotrack (suntracking). 
  • SM batteries:  Battery #8, formerly known as #7, is still disconnected in slot #8 for troubleshooting; all other batteries (7) are in “Partial Charge” mode.  
  • FGB batteries:  Battery #4 is off (capacity restoration mode, ROM); all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22).

Propulsion System:

  • Total propellant load available:3637 kg (8018 lb) as of 1/1/04 [SM(755) + FGB(2530) + Progress M(352) + Progress M-1(0)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH  XVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -9.4 deg, roll: 0 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at MBS PDGF #1/LEE B, with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 7:52am EST [= epoch]):

  • Mean altitude — 369.5 km
  • Apogee — 373.9 km
  • Perigee — 365.1 km
  • Period — 91.9 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.0006554
  • Orbits per 24-hr. day — 15.66
  • Mean altitude loss in last 24 hours — 120
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 29217

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.